Ammonia NH3 chemically reacts with oxygen gas O2 to produce nitric oxide NO and water H2O

.

What mass of oxygen gas is consumed by the reaction of 9.89 g of ammonia?


1
Expert's answer
2021-03-03T01:33:04-0500

"2NH_3+\dfrac{5}{2}O_2\longrightarrow 2NO+3H_2O"

Mass of NH3 reacted = 9.89 g

Moles of NH3 = "\dfrac{9.89}{17}=0.581\space mol"

According to balanced chemical reaction

2 moles of NH3 react with 2.5 moles of O2

"\therefore0.581\space mol" of NH3 reacts with = "\dfrac{2.5}{2}\times0.581\space mol" of O2

"\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space=0.727\space mol" of O2

Mass of O2 consumed = "0.727\times32=23.27\space g"